The measure of a right triangle is 74 degrees. I have already figured out the measures of the other two angles which are 90 degrees and 16 degrees. Now the 2nd part of the question asks "what are the measures of each of the corresponding exterior angles? Explain your answer."

Answer by fcabanski(1391)

When speaking of exterior angles, we mean one exterior angle at each vertex. Interior angles and exterior angles of a triangle are supplementary. Or, an exterior angle = the sum of the measures of the opposite interior angles.


At the right angle an exterior angle = 74+16 = 90 degrees. (180-90=90 is the second way to determine it.)


At the 74 degree angle an exterior angle is 90+16 = 106 degrees.


At the 16 degree angle an exterior angle is 90+74=164 degrees.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)